OpenStudy (anonymous):

1) |3k + 4| = 8? 2) |3x + 8| = 19? 3) |7x + 1| < 13? 4) |3k + 4| = 8? 5) |7x + 1| < 13?

OpenStudy (alfie):

There you have the first equality, you can solve the other two by using the very same method. Regarding the inequalities, you will have to make a graph and find the solutions.

OpenStudy (turingtest):

I will solve the first for you, you should be able to do the rest (they are all done the same way)\[\left| 3k+4 \right|=19\]means that either 3k+4=19 or 3k+4=-19 3k=15 or 3k=-23 k=5 or k=-23/3 so k={-23/3,5}

OpenStudy (anonymous):

If |x| = +x if x >0 = -x if x < 0